OpenStudy (anonymous):

Translate to an equation: f $0.30 folders and n $1.50 notebooks total $12

OpenStudy (anonymous):

.30f+1.50n=12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you need solve for n and f?

OpenStudy (anonymous):

Yes. I have to graph the equation and use the graph to determine three different combinations of folders and notebooks that total $12

OpenStudy (anonymous):

Ok, give me a minute I have to write on my paper first

OpenStudy (anonymous):

subtitution: f=12-n 0.3(12-n)+1.5n=12 3.6-0.3n+1.5n=12 1.2n= 12- 3.6 n=8.4/1.2 notebooks = 7 subtitution notebooks = 7 find folders 0.3f+1.5(7)=12 0.3f +10.5 =12 0.3 f=12-10.5 f= 1.5/0.3 folders = 5

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

Thanks very much... How do I give a medal

OpenStudy (anonymous):

next to your name said Give Medal you just click